### v1

# Flo: A data-driven limited-area storm surge model

Abstract
We present Flo, a data-driven storm surge model, covering the North Sea, Norwegian Sea and Barents Sea. The model is built using the Anemoi framework for creating machine learning weather forecasting systems, developed by the European Centre for Medium-Range Weather Forecasts and partners. The model is based on a graph neural network, and is capable of simulating water level due to storm surge at a horizontal resolution of 4 km and a temporal resolution of 1 hour with a quality comparable to the numerical model on which it was trained. The model was trained using a data set consisting of 43 years of atmospheric data from the 3-km Norwegian Reanalysis hindcast for mean sea level pressure and winds, and the NORA-Surge hindcast for storm surge. Evaluation was done by comparing results from hindcast runs of the Flo model against independent observations of more than 90 water level gauges along the European coast, and against the NORA-Surge hindcast. The evaluation shows that Flo produces hindcasts with accuracy similar to the NORA-Surge hindcast. Since no data assimilation was applied in the NORA-Surge hindcast used as training data, the Flo model is not expected to outperform the numerical model when compared to observations. The current work takes an important step transforming storm surge forecasting from being based purely on numerical physics-based models, to taking advantage of recent advancements in machine learning. This does not represent a large step forward with regards to improving the forecast skill, but forms a foundation for future training of a storm surge model that offers more flexibility for taking observations into account .
Keywords Water level $\cdot$ Storm surge $\cdot$ Machine learning $\cdot$ Graph neural network
1 Introduction
Neural networks (NN) have been increasingly applied to storm surge prediction and post-processing, particularly for the correction of systematic model biases (e.g., tedesco23; Zhu2025PINN; Hermans2025Europe; Naeini2025HDNN). In addition, recent advances utilizing machine learning (ML) techniques and data-driven models (DDM) within the field of meteorology have resulted in a new generation of weather forecasting models based on graph neural networks (GNNs) that are efficient and lightweight to run, and in many cases surpass the forecast quality of their numerical weather prediction (NWP) model counterparts (keisler2022; bi2023; chen2023; kurth2023; Ben_Bouall_gue_2024; lang2024; nipen2025). The latter is a direct result of the fact that many of these DDMs are trained on reanalyses as training data sets. A reanalysis is largely influenced by the assimilation of observations, and hence, a DDM trained on such data, is not limited by the discretizations of the governing equations of motion which are the foundation for most NWP models. Building on this line of research, we develop a limited-area model (LAM) for predicting water levels due to storm surge based on a GNN using the Anemoi framework for creating machine learning weather forecasting systems (lang2024).
Storm surge is the atmospheric contribution to the sea surface height (SSH) due to the inverse barometer effect (wunsch1997) and the horizontal displacement of water masses by wind stress acting on the water surface (kristensen2024). Together with the periodic and highly predictable astronomical tides, storm surge is the main contribution to the total water level observed along coastlines. Due to its direct dependence on the actual atmospheric conditions, it has to be forecasted on a day-to-day basis using operational storm surge models (e.g zijl:etal:2013; kristensen22). Extreme water levels that can pose a risk to life and property are most often cause by storm surge, often in combination with high tides.Storm surge prediction is therefore aimed at skillfully describing extreme values. However, areas along the coastal zone have different tolerance thresholds for water level, and what is considered an extreme in terms of impact on infrastructure varies geographically (fang2014). Storm surge is a barotropic process, i.e., depth-integrated shallow water dynamics apply, acting on the entire water column. Storm surge can therefore be accurately simulated using barotropic two-dimensional ocean models. Due to the good quality of the present-day weather models and as a consequence of its critical dependence on the atmospheric conditions, storm surge is highly predictable at synoptic time scales.
In addition to operational forecasting, storm surge models are also used to provide data sets of historical storm surge, i.e. hindcasts. Observation data sets consisting of data from water level gauges are by many considered as the preferred product for knowledge about historical storm surge. However, these data sets often have poor coverage both geographically and temporally, therefore models fulfill a crucial role to provide spatio-temporal coverage.
Hindcasts are used for various purposes like e.g. calculating extreme and return values (kristensen2024) and defining safety zones for infrastructure located in the coastal zone. Recently, hindcasts and reanalyses are increasingly being utilized as training data sets for DDMs (e.g. graphcast2023; lang2024; nipen2025). The present work make use of the NORA-Surge hindcast (kristensen2024) as training data in the training of a DDM for storm surge modeling, the Flo model. In the current study, we train the Flo model on hindcast data, and evaluate it by running it in hindcast mode. The model can also be applied in forecast mode, even if not demonstrated here.
The article is structured as follows. Section 2 describe the model setup, including the training data and the model architecture of the Anemoi GNN model. The training and inference of the model is outlined in Section 3. Further, Section 4 contain details about the observation data set used for the model evaluation, and the evaluation of the model performance compared against both the NORA-Surge hindcast, and observations. Section 5 provide a summary and some concluding remarks.
2 Model setup
2.1 Training data
Variable Source Type
SSH (storm surge) NORA-Surge Prognostic
10 meter wind (U10m and V10m) NORA3 Forcing
MSLP NORA3 Forcing
Bathymetry NORA-Surge Forcing (static)
Land-sea mask NORA-Surge Forcing (static)
Coriolis parameter NORA-Surge Forcing (static)
Solar insolation Anemoi Forcing (static)
Sine of latitude Anemoi Forcing (static)
Sine of longitude Anemoi Forcing (static)
Sine of local time Anemoi Forcing (static)
Sine of Julian day Anemoi Forcing (static)
Cosine of latitude Anemoi Forcing (static)
Cosine of longitude Anemoi Forcing (static)
Cosine of local time Anemoi Forcing (static)
Cosine of Julian day Anemoi Forcing (static)
Table 1: List of the variables used in the training of Flo, including their source and type (i.e. if they are forcing or prognostic).
The Flo model is built using the Anemoi framework (lang2024), which is complete software suite, written mainly in python, that handles all the steps of creating machine learning weather and ocean models in a user friendly way. During the present work, we have used the anemoi-datasets for preparing and converting the training data set, anemoi-training for the training of the model, and finally anemoi-inference for running the the hindcast inference runs . It is a LAM configuration in Anemoi, which means it uses training data for a regional domain covering the North Sea, Norwegian Sea and Barents Sea (see Figure 2). The data used for training consists of two datasets: 1) the NORA3 atmospheric hindcast (haakenstad21nora3) and 2) the NORA-Surge storm surge hindcast (kristensen2024). Variables used in the training, see Table 1, are the SSH from the NORA-Surge hindcast and as forcing variables the 10 meter winds as x- and y-components (U10m and V10m) and mean sea level pressure (MSLP) are taken from NORA3. All variables are interpolated onto a common grid with a horizontal resolution of 4 km. The temporal resolution of the training data is 1 hour. We note that the NORA-Surge hindcast was forced by wind stress calculated from NORA3 winds based on the Charnock relation (charnock:1955), while in the present work, we have opted to use the wind directly in the training. Other forcing variables include bathymetry, sea-land mask, and the Coriolis parameter. In addition, other standard ML-forcings added by the Anemoi-framework include insolation, and the sine and cosine of latitude, longitude, local time and Julian day.
A total of 43 years of data are available from the NORA3 and NORA-Surge hindcasts.The model training uses the 16 years time period 1990-2005 (inclusively), while the validation data set, used for training validation metrics, consists of the four years 2006-2009. The remaining 13 years 2010-2022 were reserved for evaluation of the model and case studies. The selected periods, and their lengths, are in part chosen to match the time periods with large amounts of observation data to compare with, and to enable us to complete the training within reasonable time on the available compute resources.
Figure 1a shows the statistical distribution of storm surge for each year of the NORA-Surge data set (all grid points), with the training, validation and evaluation periods marked with a blue, green and yellow background, respectively. In Figure 1b we show the summarized distribution of storm surge for the training, validation and evaluation periods. The data set contain, for each time step, a total number of 593 920593\,920 grid points, whereas 436 047436\,047 of them (approximately $73\%$) are ocean points. It is clearly evident in Figure 1 that the water level due to storm surge is very close to zero most of the time. However, the socio-economic impact of storm surge, including consequences to life and property, becomes increasingly important as the amplitude of the storm surge increase, i.e., it is a case of extreme value prediction. Figure 1 show that even in a 43 year long data set, extreme values are very rare.
Usually, LAM-models implement lateral boundary forcings using a lower-resolution model surrounding the regional domain (e.g., see oskarsson2023; adamov2025; wijnands2025 which uses coarser global models as boundary forcing datasets). In an operational setup, a data-driven LAM model requires boundary forcings at each forecast time step, which may be provided by an external numerical model. In the current work we have taken a simpler approach, where we have reduced the size of the prediction domain by five grid points, and used the remaining cells as boundary forcings. This means that the state of boundary domain is provided by the same dataset as the interior domain, and that the resolution is the same. This setup may be extended to facilitate an operational setup, as further discussed in 5.
Figure 1: Histograms showing the distribution of storm surge in the entire NORA-Surge training data set, for all times and all grid points. Panel (a) show a 2D year-by-year comparison of storm surge distribution (values on the color scale are logarithmic), whereas panel (b) show combined distribution for the relevant training, validation and evaluation periods, also on log scale. The blue, green and yellow background colors in panel (a) depict the training, validation and evaluation periods, respectively. In addition, the horizontal dashed red line mark the year 2013 that contains the storm Xaver.
2.2 Model architecture
Flo is based on an encoder-processor-decoder architecture, similar to the works of graphcast2023; nipen2025; lang2024. The encoder transforms the input state, which consists of surface variables and weather forcings relevant to storm surge prediction (see Table 1), into a lower-dimensional latent space (see Figure 3a). Within this latent space, referred to as the processor mesh (see Figure 4), the processor advances the state forward in time. Finally, the decoder maps the predicted latent state back to the original data grid (as shown in Figure 3b), providing the desired storm surge variables on the input grid. Within Anemoi, we apply the graph transformer option for the model with 16 attention heads, as described in lang2024. The setup as graph transformer network allows to propagate information over longer distances across.
Figure 2: The storm surge model domain covering the North Sea, the Norwegian Sea and the Barents Sea is shown by the shaded area.
(a) Encoder
(b) Decoder
Figure 3: Example of how the mapping of the physical model grid (data grid) space is mapped into the hidden mesh via the encoder in panel (a) and back into data grid via decoder in panel (b).
(a) Processor outer
(b) Processor inner
Figure 4: Visualization of how the processor connects the nodes of the hidden mesh via edges at different refinement levels. Panel (a) show the edges between nodes at refinement level 4-6, where lower number equals longer edges. Panel (b) show the edges and connections between the nodes at refinement levels 7-10 for the area outlined by the square of dotted lines in panel (a).
The graph is a central piece of the GNN architecture, consisting of nodes in data and latent space that are connected by edges. The data nodes are created directly from the grid points of the NORA3 and NORA-Surge training data. Following lang2024 and graphcast2023, the processor mesh nodes are represented by the vertices derived from iterative refinements of an icosahedron around the globe, see Figure 4. At base refinement $N=0$ we obtain an icosahedron with $12$ vertices, $30$ edges and $20$ faces. Each increase in refinement level divides the existing faces into four smaller ones, increasing the number of vertices, and thus the number of processor mesh nodes. The number of vertices $V_{N}$, or mesh nodes, at refinement level $N$ can be expressed as:
$V_{N}=2+10\cdot 4^{N}.$ (1)
Our mesh is constructed at refinement level $10$, yielding 10 485 76210\,485\,762 global processor mesh nodes. However, as our model uses a LAM setup, we remove any processor mesh nodes outside the region of interest, yielding 211 525211\,525 processor mesh nodes, roughly one third as many processor nodes as grid nodes.
The encoder maps information from data nodes to hidden nodes through the edges, see Figure 3a. To create the edges of the encoder, we evaluate the distance between two neighboring processor nodes. This distance is the largest permitted encoder edge length. Encoder edges are created to connect each processor mesh node to all data grid nodes within this distance. Thus a single processor mesh node will obtain information from multiple data grid nodes, and data grid nodes send information to multiple processor mesh nodes. To reduce the complexity of the graph, the maximum number of data nodes that a mesh node may be connected to is $64$. Thus, if the number of edges reaches $64$ during edge creation, the algorithm will stop, which may leave unconnected nodes called orphans. At refinement level $10$ we do not observe such unconnected nodes.
Similarly to the creation of the processor mesh nodes described above, the processor edges are directly obtained from the edges of the iteratively refined icosahedron (e.g. Figure 4). For a global model with refinement $N=0$, each processor node will communicate with the five neighboring processor nodes. Each increase in refinement level $N$ creates new nodes which communicate with its neighbors. The length $L$ of a processor edge at each refinement $N$ can be described as:
$L_{N}=\frac{L_{N-1}}{2},$ (2)
i.e. for each increase in refinement level the edge lengths are halved. Processor edges from all iteratively constructed refinement levels are accumulated.
For our LAM domain we remove any edges and nodes outside of the model domain (see Figure 2). This effectively removes all edges at $N=0$ as these are too long. Even if storm surge is a fast moving phenomenon that effectively travels the length of the model domain in less than 24 hours, the few scattered long edges on refinement level 1-3 only provide minimal coverage due to the orientation of the polar stereographic domain. Thus we decided to only keep processor edges at refinement levels $N\geq 4$, which further reduces the complexity of our graph.
The decoder maps information from the processor mesh in latent space back to the data grid. Our decoder connects each data grid node to the three closest processor mesh nodes by using the k-nearest neighbors algorithm (KNN). Our LAM setup uses a boundary region to provide the model with lateral boundary conditions. Since we do not want to predict the ocean state at the domain edge, we construct our decoder such that no mesh nodes are connected to data nodes in the boundary region.
3 Training and Inference
3.1 Training
We have performed two trainings during the development of the Flo model. One is a training based on 16 years (1990-2005) of data, hereafter referred to as the Flo model, while the other was based on the 4 years 1990-1993. Even if the NORA-Surge hindcast covers a longer time period, we have selected the years starting from 1990 because more observations are available from 1990 and onwards. The two training runs enable us to evaluate if we should prioritize a longer training period containing more data, or if running a shorter period with less data for more epochs provides better result given the same amount of GPU resources. Both training periods could be considered to be quite short, but based on the distribution of storm surge for the years that make up the training data set, as shown in Figure 1, it is clear that both of the selected training periods contain a distribution of storm surge that to a large extent represent the overall distribution. In addition, we suspect that the physical processes that drive the storm surge (i.e. the inverse barometer effect and the advection due to wind stress) are easily identified by the model training, and does not require a very long training period with many extremes in order to be able to simulate extremes.
The model was trained on a single Nvidia H200 GPU using Anemoi-training version 0.6.4 on local infrastructure at MET Norway. We used a learning rate of $6.25\times 10^{-4}$, and the model was run for $10^{5}$ steps with a batch size of $4$. The warmup period was set to 1 0001\,000 steps, and the minimum learning rate of the cosine learning rate schedule was $3.0\times 10^{-7}$ and the number of channels was $256$. The same setup was used for both the trainings, and results in $\textasciitilde 3$ and $\textasciitilde 12$ epochs for the long and short training periods, respectively. Each of the two trainings consumed $142$ GPU hours.
Figure 5: The training loss for every 100 training steps as a function of steps, as calculated by Anemoi-training. We show the training loss for both of the training runs, together with a smoothed line for each of the two.
The training MSE loss, as calculated by the Anemoi framework during training, for the two runs are displayed in Figure 5 on log scale. The loss start at around $5.0\times 10^{-3}$ for both trainings, and quickly drops below $10^{-3}$ during the first 1500-2000 steps. It is clear how the loss curves still have a negative trend even towards the end of the training period, suggesting that the training could benefit from running for more steps.
3.2 Inference
The evaluation of the Flo model, presented in Section 4, is based on inference of the model as a series of hindcast runs. Firstly, we perform a continuous run of the DDM, initialized from NORA-Surge at 2010-01-02 00 UTC and run for 78840 hours until 2018-12-31 00 UTC. The inference was forced by NORA3 10 meter winds and MSLP, and the SSH boundary data was taken from the NORA-Surge hindcast. The runs was done on the same H200 GPU for inference as was used in the training. We performed the hindcast runs twice, based on the training checkpoints from both the 4 and 16 years long training runs after $10^{5}$ steps.
In addition, we have conducted a series of shorter inferences of 240 hours length. Inference was run for the checkpoints after 25, 50, 75 and $100\times 10^{3}$ steps for both training runs. These were re-initialized from NORA-Surge every 10 days during the year 2013, resulting in 36 runs for each of the checkpoints, a total of 288 runs. Atmospheric and boundary forcing was the same as for the longer hindcast runs. This was done to explore at which rate the two training runs converge towards NORA-Surge.
4 Evaluation
The quality of the Flo model is evaluated for the time period 2010-2019, which has not been seen by the model during training. We present here a direct comparison with the results from the same numerical model on which Flo was trained, and a comparison with water level observations. We note that we do not expect superior performance of the DDM copared to the hindcast, contrary to many DDMs for atmospheric prediction that realize better performance than their NWP model counterparts due to training on reanalysis (e.g. graphcast2023; lang2024; Ben_Bouall_gue_2024; nipen2025). We do not expect Flo to perform better than the physical model since it is not influenced by observations, but rather a ML model emulating the physics of a numerical model. Hence, a "perfect" result would be that the DDM produces exactly the same fields as those obtained by running the numerical model, i.e. the NORA-Surge hindcast.
4.1 Observation data set
Figure 6: Positions of the observation locations used in the analysis.
The observation data set used to evaluate the quality of the Flo model consist of a collection of water level observations from 98 tidal gauge stations around the North Sea, Norwegian Sea and Barents Sea (see Figure 6). The same data set was quality controlled and used for validation by tedesco23 and kristensen2024. The observation data set is de-tided using harmonic analysis, as described by kristensen2024, and the following analyses focuses purely on the storm surge part of the water level signal.
4.2 Comparison against NORA-Surge
Figure 7: Time series of RMSD between Flo and NORA-Surge at every timestep during the inference runs (2010-2018). Panel (a) show the first 480 hours of the inference run, while panel (b) show the entire time period.
To evaluate the difference between Flo and NORA-Surge, we compute the root-mean-squared-differences (RMSD) between the two for every grid point for all time steps: Figure 7 shows that the RMSD increase during the first 24-48 hours, and stabilize between 2 and 3 cm for the remainder of the initial period (Figure 7a). This could be an effect of smoothing, which is a known behavior in many ML-models (graphcast2023, e.g.). A careful examination of the RMSD evolution in Figure 7b reveal a few key observations: Firstly, it the RMSD exhibits seasonality and secondly, peaks in RMSD increase for lead time in the inference. In addition, there is a slight positive trend in the RMSD. The comparison between the two trainings based on 4 and 16 years of data suggests that the longer training reveals slightly lower RMSD on average, but generally differs little.
Figure 8: RMSD between inference of the Flo model based on the two training runs (4 and 16 years) at different steps of the training runs, and NORA-Surge. The RMSD, in meters, is calculated as an average over all grid points at every hour for the entire 240 hours inference run for each of the 36 inference runs performed for each of the checkpoints for both trainings (as described in Section 3.2).
In addition, as mentioned in Section 3, we explore how the two different lengths of the training data set and the number of training steps affect the model performance. Figure 8 show a comparison of the average RMSD between inference of the Flo model and NORA-Surge based on the two training runs at four different checkpoints as described previously in Section 3.2. This suggests that similar quality for the DDM is obtained, as long as the runs are sufficiently long . The test in Fig. 8 indicates that the DDM based on the long training period converges faster, but we emphasize that the rate of convergence can depend on the time period and cannot be generalized.
4.3 Comparison against observations
One of the most important metrics of any storm surge model is how well it compares to observations. We compare against the observations introduced in Section 4.1. Prior to comparison, we remove the mean from both the observation and model data sets to account for differences in sea surface reference levels. The evaluation focuses on the root-mean-squared-error (RMSE) and correlation, and we compare the hindcasts (see Section 3.2) based on both the short and long training period, and the NORA-Surge, against the observations.
To explore how well the three hindcasts perform with regard to simulating storm surge of significance, we evaluate and validate the results after masking time periods with absolute values of the observed storm surge less than 10 and 30 cm. Since the different stations have different distributions of storm surge, and it is common that higher absolute values of storm surge result in higher RMSE values, we also include the average RMSE scaled by the standard deviation of the observed storm surge for each station. The results for all stations combined is displayed in Table 2, and owing to the fact that our main geographical area of interest is the coast of Norway, we include the results calculated for the Norwegian stations only in Table 3.
[…]
NORA-Surge All data 12.6 0.55 0.85
Masked $\pm 10cm$ 13.4 0.58 0.90
Masked $\pm 30cm$ 16.7 0.71 0.94
Train 16 years All data 12.4 0.54 0.86
Masked $\pm 10cm$ 13.2 0.57 0.90
Masked $\pm 30cm$ 16.6 0.70 0.94
Train 4 years All data 13.1 0.58 0.84
Masked $\pm 10cm$ 13.9 0.61 0.89
Masked $\pm 30cm$ 17.4 0.75 0.94
Table 2: Averaged statistics for all 98 stations over the time period 2010-2018. The rows labeled "All data" contain all valid data from the comparison against observations. The rows labeled "Masked" contains only the data for when the absolute values of the observed storm surge is larger than the given value. I.e. all data for when the absolute value of the observations are less is masked out. This is done to explore how well the different models handle larger values of storm surge (extremes).
[…]
NORA-Surge All data 8.4 0.53 0.87
Masked $\pm 10cm$ 8.8 0.56 0.92
Masked $\pm 30cm$ 10.3 0.65 0.97
Train 16 years All data 8.1 0.51 0.88
Masked $\pm 10cm$ 8.5 0.54 0.93
Masked $\pm 30cm$ 10.0 0.63 0.97
Train 4 years All data 9.2 0.58 0.85
Masked $\pm 10cm$ 9.6 0.61 0.91
Masked $\pm 30cm$ 11.2 0.71 0.97
Table 3: Same as Table 2, but for the 24 Norwegian stations only.
(a) NORA-Surge
(b) Train 16 years
(c) Train 4 years
Figure 9: Heatmap scatter plots comparing the hindcast and observations for the Norwegian stations. Note the log scale for count. The ellipsoid displayed show the area of 5 standard deviations along, and normal to, the diagonal, centered at the point of maximum distribution. Lengths of the major and minor axis are displayed in each plot, and provide a measure of the spread around the centroid along the diagonal.
The focus on the Norwegian stations is continued in Figure 9, where we display the heatmap scatter plots that compare the hindcast and observations for the Norwegian stations. In addition, we display an ellipsoid that show the area of 5 standard deviations along and normal to the diagonal, centered at the point of maximum distribution. Lengths of the major and minor axis are displayed in each plot, and provide a measure of the spread around the centroid along the diagonal. Comparing the length of the minor axes of the ellipsoids for the three hindcast, we demonstrate that the Flo model based on the 16 year training yields the lowest error, i.e. spread around the diagonal, with the NORA-Surge and the short training at slightly higher errors. This reflects the RMSE shown in Table3.
4.4 Case study of Storm Xaver
Figure 10: Comparison of storm surge water level between Flo and NORA-Surge at the peak of the storm Xaver in the German Bight at 05 UTC on the 6th of December 2013. Left panel is NORA-Surge, middle panel is Flo and the right panel is the difference between the two.
Figure 11: Time series comparison of the observed storm surge water level for four selected stations compared to the data-driven Flo model and NORA-Surge during the storm Xaver in December 2013. The stations are, from top to bottom, Cuxhaven (GER), Esbjerg (DEN), Oslo (NOR) and Viker (NOR). RMSE and correlation for the depicted time period for each of the model runs is given in the legend for each station.
Storm Xaver moved slowly from west to east across Northern Europe during 5-7 December 2013 (deutschlander:etal:2013; kristensen2024; kristensen2025). The strong northerly winds over the North Sea resulted initially in a convergence of water along the eastern coast of the United Kingdom due to Ekman transport of water towards the coast. When the wind direction started to back left, this water started to move anti-clockwise along the North Sea coastline as a free Kelvin wave. The combined Kelvin wave with local amplifications due to Ekman transport by the wind, and the inverse barometer effect resulted in a storm surge among the top five highest recordings over the last 100 years in the German Bight according to deutschlander:etal:2013.
As the low pressure system moved further east, and the wind direction continued backing further left, the storm surge signal moved further through the North Sea up along the western coast of Denmark and into the Skagerrak area as a Kelvin wave (see kristensen2024).
The storm Xaver, due to its characteristics as an extreme and special event, has been thoroughly studied by e.g. staneva17 and kristensen2024; kristensen2025. Figure 10 show the storm surge in the North Sea from the Flo model compared to NORA-Surge at the peak of the storm surge in the German Bight during storm Xaver. The fields from the two hindcasts are visually very similar, and the difference-plot confirm that even the largest differences are less than 20 cm. The general impression of the difference is that the Flo hindcast has lower amplitudes than NORA-Surge most places, with a few exceptions. An area that stands out from the rest is the Skagerrak area between Norway, Denmark and Sweden, where the Flo hindcast has higher amplitudes than NORA-Surge. This is generally an area where NORA-Surge, and the operational model based on the same model configuration, struggle (see kristensen22; tedesco23; kristensen2024). The time series in Figure 11 show a comparison between observations and the three hindcasts for four selected stations from Cuxhaven (GER) in the south, up to Oslo (NOR) in the north. This clearly show how the Flo model based on the 16 year training period outperform the NORA-Surge, and the inference based on the shorter training period, for the two Norwegian stations located in the Skagerrak area for the storm Xaver, both with regards to RMSE and correlation.
This suggest that, contrary to some DDMs for the atmosphere and weather prediction, that emphasize their shortcoming when it comes to predicting extremes (e.g. nipen2025), the Flo model is able to reproduce an extreme event like storm Xaver with similar, or better, accuracy than the NORA-Surge based on a numerical model. This is likely an effect of smoothing of the model fields, so the effect of double penalty is reduced. This is a common and well known behavior in GNN based ML models as shown by e.g. graphcast2023 and nipen2025.
5 Summary and conclusions
We present Flo, a data driven storm surge model capable of simulating reliable storm surge predictions for the limited area model domain covering the North Sea, the Norwegian Sea and the Barents Sea. Flo produces storm surge simulations at a horizontal resolution of 4 km, and a temporal resolution of 1 hour. The model is built using the Anemoi framework (lang2024), and utilizes a graph neural network to predict storm surge. The training of Flo is based on the NORA3 hindcast (haakenstad21nora3; haakenstad22) for atmospheric forcing (MSLP and 10 meter winds), and the NORA-Surge hindcast (kristensen2024) for storm surge. Both hindcasts cover the 43 year period from 1979 to 2022, and we utilize the 20 year period from 1990 to 2009 for training and validation. The years 2010-2022 was reserved for evaluation. Hindcasts covering the period from 2010 to 2018 created by Flo have been evaluated qualitatively and validated against both the NORA-Surge hindcast and a data set of observations of storm surge from 98 water level stations. The evaluation demonstrate that Flo has similar – or slightly better – skill than its numerical counterpart, here represented by the NORA-Surge hindcast. When averaging over the entire time period of the Flo hindcast, and all stations, NORA-Surge yields an RMSE of 12.6 cm, while Flo yields an RSME of 12.4 cm. This improvement holds even when masking out the time periods with observed absolute values of storm surge of less than 10 and 30 cm. We note that this improvement is likely an effect of smoothing of the model fields, reducing the effect of double penalty.
The case study of the extreme winter storm Xaver of December 2013 shows that Flo is capable to predict an extreme and rare storm surge event with similar quality as NORA-Surge. For some stations, Flo delivers better predictions.
We acknowledge the fact that developing and training a DDM, like Flo, is a costly process with regards to computing resources, and is critically dependent on having an existing hindcast or reanalysis for training. With this in mind, a relevant question is whether the marginal improvements that this model currently realizes can be justified. We acknowledge that the full potential of transitioning from a purely numerical model to a data-driven machine learning model lies not in the here presented work, but the potential for future developments: As a next step, observations could be included in the model training to improve the predictions. By doing this, the DDM can develop the ability to predict storm surge more accurately than the numerical model, since it is not limited by the discretizations of the governing equations of motion, but is able to find statistical relations in the training data which includes information from observations. In our case, instead of running a costly reanalysis using a numerical model, the inclusion of information from observations may be performed by correcting the hindcast using primitive methods, like e.g. optimal interpolation. Future work could include further training on operational forecast data from the numerical model used for storm surge (kristensen22), and possibly involve training on ensemble model data to prepare the Flo model for operational forecasting runs and application in decision-making from impact-based forecasting services.